OpenStudy (anonymous):

This is a calculus 2 question regarding population models. 1000(dp/dt)=p(100-p) The initial population is 200 individuals. Find p(t)

OpenStudy (anonymous):

looks like a calc 1 prob to me. can you not take the integral in that form?

OpenStudy (anonymous):

1000dp/(p(100-p)) = dt dp/(p(100-p)) -> A/p + B/(100-p) = 1000/(p(100-p)) A = 10 , B = 10 (10/p + 10/(100-p))dp = dt integral of both sides .. 10ln(p) - 10ln(100-p) = t + C 10ln(p/(100-p)) = t + C ln(p/(100-p)) = t/10 + C (C is constant so i can divide it by 10 and still call it "C" .. if you want call it now C2) p/(100-p) = e^(t/10 + C) p/(100-p) is like -1 + 100/(100-p) so -1 + 100/(100-p) = e^(t/10) * e^C 100/(100-p) = Ce^(t/10) + 1 100 = 100Ce^(t/10) + 100 - p(Ce^(t/10) + 1) 100Ce^(t/10) = p(Ce^(t/10) + 1) p = 100Ce^(t/10) / (Ce^(t/10) + 1) and you can change it to : p = 100 - 100/(Ce^(t/10) + 1) where C = -2 since p(0) = 200